OpenStudy (destinykiara99):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f of x equals eight divided by x and g of x equals eight divided by x

OpenStudy (shawn):

f(x) = 8/x g(x) = 8/x

OpenStudy (destinykiara99):

yes

OpenStudy (shawn):

f(g(x)) = 8/(8/x) = 8 * x/8 = x

OpenStudy (destinykiara99):

Would you just do the same to g(x) then?

OpenStudy (shawn):

yes, do the same g(f(x))

OpenStudy (destinykiara99):

g(f(x))= 8/(8/x) =8*x/8 =x it's all exactly the same right?

OpenStudy (princeharryyy):

yes! it is.